OpenStudy (anonymous):

stepswise solution product of 4 consequtive natural nos. is 5040 find them

OpenStudy (raden):

the factors of 5040 = 2^4 * 3^2 * 5 * 7 = 2^3 * 2 * 3^2 * 5 * 7 = 7 * 8 * 9 * 10